# Overestimation of melting temperatures calculated by first-principles molecular dynamics simulations

### v1

@@ II.1.2 Construction of MD simulations
In this study, the internal energy $U$ must be evaluated over a wide range of $T$ to cover from the solid to liquid states. To ensure that equilibrium was established at each $U$, an adiabatic MD run was performed for a fixed $E$ and repeated by changing the total energy. In adiabatic MD simulations, the control parameter is the initial temperature $T_{\rm in}$, at which the initial velocities of atoms are given through Eq. (6). In each single MD run, the simulation was continued until equilibrium was established. It is noted that there are two kinds of equilibrium conditions Shirai (2020). The one is the equilibrium condition with respect to temperature (thermal equilibrium). The equilibrium temperature is determined by
@@
Here, the brackets denote the particle average, whereas the bar indicates the time average. Thermal equilibrium is reached when the time average of $E_{\rm K}(t)$ converges. In the following, $T$ in all figures indicates this equilibrium temperature. The time required to reach this condition is referred to as the thermal relaxation time, $\tau_{T}$. For most of solids, $\tau_{T}$ is quite short, on the order of 0.1 ps. For liquids, $\tau_{T}$ becomes longer but is still less than 1 ps. Thus, in most cases the relaxation to equilibrium is controlled by the other condition, namely, the mechanical (structural) relaxation. The time required to this relaxation is referred to as the structural relaxation time, $\tau_{M}$. The time evolution of the particle-averaged displacements,
$\langle\overline{\delta R_{j}(t)^{2}}\rangle_{j}=\frac{1}{Nt_{\rm SM}}\sum_{j}^{N}\int\left(R_{j}(t-t_{0})-R_{j}(t_{0})\right)^{2}t_{0},$ (9)
@@
is used to determine whether structural equilibrium is reached, where $t_{\rm SM}$ is the simulation time. For solids, it is considered an equilibrium state when $\langle\overline{\delta R_{j}(t)^{2}}\rangle_{j}$ is constant with respect to $t$. In contrast, for liquids, it is considered an equilibrium state when $\langle\overline{\delta R_{j}(t)^{2}}\rangle_{j}$ shows a linear dependence on $t$ over the entire simulation time $t_{\rm SM}$. For the latter case, the slope, $D^{\prime}=\langle\overline{\delta R_{j}(t)^{2}}\rangle_{j}/t$, gives the diffusion coefficient. Although the true diffusion coefficient, $D$, is given by $D=(1/6)\lim_{t\rightarrow\infty}\langle\overline{\delta R_{j}(t)^{2}}\rangle_{j}/t$, $D^{\prime}$ is used for the diffusion coefficient in the following. For the liquids treated in this study, $\tau_{M}$ was on the order of a few ps. Examples of the time development of $E_{\rm K}$ and $\langle\overline{\delta R_{j}(t)^{2}}\rangle_{j}$ are provided in the Supplemental Material to Ref. Shirai et al. (2023).
II.2 Calculation conditions
@@
While several codes were used for the present FP-MD simulations, the majority of calculations were performed by Phase/0 Yamasaki et al. (2019). Some were performed by a code written in-house (Osaka2k) or the VASP code VAS . All codes are pseudopotential methods using planewave expansion. The MD simulations employed time steps ranging from 0.72 to 1.2 fs and the total simulation time, $t_{\rm SM}$, varied from 2 to 10 ps, depending on the relaxation time. The volume of the crystal, $V$, was fixed at the experimental value, although normally the $V$ of the corresponding liquid phase is slightly larger than that of the crystal. The effect of volume change on $T_{m}$ is examined in subsection III.4.
III Results and discussion
III.1 Melting temperature
@@
Melting simulations for various materials were performed by FP-MD simulation by scanning a wide range $T$ covering both the solid and liquid states. Table 1 summarizes the calculated $T_{m}$ together with experimental values for comparison. The potential type is indicated by NC (norm-conserving), US (ultrasoft), PAW (projector augmented waves) Martin (2004), along with the type of electron correlation energy functional (LDA or GGA). The cutoff energy $E_{\rm cut}$ for the planewave expansion is given in units of eV. In MD simulations, melting occurs over a finite range of $T$, as shown later. Here this $T$ range is referred to as the transition region. The calculated value of $T_{m}$ is then defined as the mid point of the transition region. The width, $\Delta T_{m}$, of the transition region is expected to approach zero as the size of supercell is increased. In the table, the deviation of the calculated value from the experimental value is indicated by the ratio, $r$, of the former to the latter. In all the cases studied, $r>1$ was obtained, meaning that the calculated values were always overestimated. It is our surprise to obtain large values of $r$ more than 2, despite the initial confidence in the DFT calcluations. Large overestimations of $r>2$ are found when the material contains oxygen atoms. Only for Si, the calculated value is in a good agreement with experiment.
@@
In the table, the melting enthalpy $H_{m}$ is also listed in units of eV/atom. $H_{m}$ is obtained from the abrupt change in $E_{\rm st}$. In experiment, the change $H_{m}$ appears at a discrete temperature $T_{m}$, whereas, as noted, melting occurred over a finite width $\Delta T_{m}$ in the present MD simulations. Hence, the accuracy of $H_{m}$ depends on the sharpness of the transition region. Small cell sizes give less reliable $H_{m}$ values. The accuracy of $H_{m}$ highly depends on the accuracy of $T_{m}$. These two quantities are related each other by
@@
where $\Delta S_{m}$ is the entropy of melting. This relationship says that an error of 0.1 eV in $H_{m}$ causes an error of 1000K in $T_{m}$, provided $\Delta S_{m}/k_{\rm B}=1$. Because $\Delta S_{m}$ in many materials is of this order of magnitude, the accuracy of $H_{m}$ directly controls the accuracy of $T_{m}$. Again the calculation for Si shows a good agreement in $H_{m}$ with experiment. This confirms the internal consistency between calculated $T_{m}$ and $H_{m}$ for Si. The phonon energy, $E_{\rm ph}$, and thermal expansion energy, $E_{\rm te}$, can contribute $H_{m}$. However, the previous studies Shirai et al. (2022, 2023) showed that these two contributions can generally be ignored, unless there is a large difference in thermal expansion between the solid and liquid phases. Because the melting temperature of most materials is high, the specific heat of the phonon contribution near $T_{m}$ becomes already the classical limit of the Dulong-Petit law, $C_{v}=3k_{\rm B}$, and thus there is no reason that an abrupt jump in $E_{\rm ph}$ occurs at $T_{m}$.
@@
$\alpha$-quartz (SiO2) 1700 4500 2.64 0.033 0.13 72 US+GGA 408 Phase
Glycerol (C3H8O3) 291 635 2.18 0.013 0.010 56 US+GGA 408 Phase
@@
Table 1: Comparison of calculated $T_{m}$ to the experimental value for various materials. $r$ indicates the ratio of the calculated $T_{m}$ to the experimental value. The melting enthalpy $H_{m}$ is also listed. Size indicates the size of the supercell used in MD simulations. The last block indicates the used potentials. $E_{\rm cut}$ for planewave expansion is given in eV units. $\Gamma$-point $k$ sampling is used in all the listed data. Experimental data are from CRC (2011) with the exception of the data for quartz Richet et al. (1982).
@@
Several causes for the error in the calculated $T_{m}$ are investigated. One important cause is the use of periodic boundary conditions. An effect of the periodic boundary conditions is elimination of the surface effects. Because melting first occurs at the surface of a material, it can be argued that elimination of the surface makes melting less likely, thus increasing $T_{m}$. This effect could contribute to the overestimation of $T_{m}$. However, based on the good agreement between the calculated and experimental $T_{m}$ values for Si, this cannot be the primary reason for the large overestimation of $T_{m}$ unless a very large surface energy is involved. The introduction of defects could have a similar effect on melting. However, in our experience for quartz Shirai et al. (2023), the introduction of a defect in a supercell did not affect $T_{m}$ within the numerical accuracy.
@@
Another effect of the periodic boundary conditions is the creation of spurious reflections of atom movements, which yields an artificial barrier to such movements. For a given size of supercell, $L$, all long-wavelength phonons (those with wavelength $\lambda>L$) are eliminated. In a previous paper Shirai et al. (2023), by using a crude model, it is shown that this artificial barrier, $V_{0}$, is reduced only in a scale $1/L^{2}$, while the computational demands increase as $L^{3}$. The value of $V_{0}$ can be of the order of a few tens of eV with a cell size of $2\times 2\times 2$ for $\alpha$-quartz. However, to test this scaling, real calculations is needed by increasing the cell size. In practice, extending the size of supercells beyond those listed in Table 1 is not feasible. For a 512-atoms simulation, one month computation was required for a single temperature in our cluster machines (typically in an 8-core parallel configuration). Hence, the effect of supercell size can realistically only be examined by reducing the supercell size, even though simulations become unrealistic. In the following, the effect of supercell size is examined by taking one example from each of the categories of semiconductors, metals, and oxide compounds. All the detailed data listed in Table 1 are seen in the Supplemental Materials.
III.2 Silicon
@@
The first example is Si. The cell sizes of $N=$8, 64, 216, and 512 atoms were examined. In Fig. 2, the structural energy, $E_{\rm st}$, and diffusion constant, $D^{\prime}$, are plotted as functions of $T$. Note that $E_{\rm st}$ is normalized by the number of atoms in the cell and is in units of K. This is convenient because the specific heat is immediately obtained in units of $k_{\rm B}$, making it easier to compare the obtained value to the classical limit $3k_{\rm B}$. This convention is used throughout this paper. In this figure, the data points are connected by lines to show the sequence of changing $T_{\rm in}$. As shown in Fig. 2, $E_{\rm st}$ is essentially constant with changes in $T$ up to $T=1000$ K, indicating that the crystal potential is well described by the harmonic approximation. Finite values of $D^{\prime}$ appear approximately $T=2000$ K, confirming the onset of melting. In accordance with the variation in $D^{\prime}$, $E_{\rm st}$ exhibits an abrupt increase.
@@
Figure 2: The diffusion coefficient, $D^{\prime}$, and structural energy, $E_{\rm st}$, of Si calculated by NC+GGA. The cell sizes ($k$ meshes) are $N=8$ ($4^{3}$), 64 ($2^{3}$), 216 (1), and 512 (1). $E_{\rm cut}=204$ eV. The experimental value for $T_{m}$ is indicated by the dashed line. For each $N$, the energy origin is set to the ground state energy of that cell.
@@
Sigmoidal behaviors with negative slopes are observed in both $E_{\rm st}$ and $D^{\prime}$ around $T_{m}$. The negative slope in $E_{\rm st}$ against $T$ means a negative specific heat, which is thermodynamically unstable Lynden-Bell (1999); Landsberg and Pearić (1987); Michaelian and Santamaria-Holek (2007). The negative slope in $D^{\prime}$ against $T$ also means dynamical instability. This behavior was pointed out previously as the effect of finite-size cells Shirai et al. (2022, 2023). In a finite-size MD simulation with $N$ atoms, the kinetic energy has a fluctuation of $\Delta T/T=1/\sqrt{N}$. When the crystal melts, the energy distribution between the vibrational motions (in the solid portion) and translational motions (in the liquid portion) becomes unbalanced due to this energy fluctuation. At temperatures slightly below $T_{m}$, those atoms with sufficiently high kinetic energy (higher than $k_{\rm B}T_{m}$) begin to convert their vibrational motions to diffusing motions, decreasing the average kinetic energy $\langle E_{K}\rangle$ and increasing diffusion. In contrast, when this low-energy part of tge distribution are overpopulated, the system temperature is lowered. This makes these low-energy diffusing atoms being trapped in the crystal potential, leading to slight heating and a decrease in $D^{\prime}$. In this manner, an oscillatory behavior appears around $T_{m}$. In experiment, this sigmoidal behavior has been observed in clusters containing small numbers of atoms Schmidt et al. (2001). The transition region is narrowed as the size of the cell increases. On the limit $N\rightarrow\infty$, $\Delta T_{m}\rightarrow 0$ is reasonably expected. However, it is noticeable that $\Delta T_{m}$ of 300 K remains even for the cell size of 512 atoms. The convergence of the energy fluctuation is slow.
@@
For Si, even though the convergence for the width $\Delta T_{m}$ is very slow, the calculated $T_{m}$, which is obtained by the mid point in the transition region, already agreed with the experimental value, when a cell size of $N=64$. Even a small cell of $N=8$ is not bad, aside from the $\Delta T_{m}$ value. From the viewpoint of construction of FP pseudopotential, Si is the easiest element Bachelet et al. (1982); Troullier and Martins (1991). The potential is not deep and there are no semi-core levels. This may be the reason why the good agreement is obtained for $T_{m}$.
@@
For the liquid state, a decrease in $E_{\rm st}$ against $T$ can be seen in Fig. 2. This negative $T$ dependence is unrealistic. In contrast, the total energy, $E_{\rm tot}$, is an increasing function of $T$, as should be. A full description of the $T$ dependence of $E_{\rm tot}$ will be given elsewhere. Here, we only mention that this decreasing behavior of $E_{\rm st}$ is a consequence of applying Bose-Einstein statistics to the “phonons” of the liquid state. The decrease in $E_{\rm st}$ means that an overly large $E_{\rm ph}$ value was subtracted from $\overline{E_{\rm gs}(t)}$ in Eq. (7). This $T$ range ($T>T_{m}$) is already the classical regime, in which the Dulong-Petit law is expected to hold, if phonon theory is valid. All “phonon” modes are fully activated, which means that the thermal energy $(1/2)k_{\rm B}T$ is assigned to each mode $q$ of phonons in Eq. (5), for which the independent particle picture for phonons is assumed. In experiment, the $C_{v}$ of liquids decreases from approximately $3k_{\rm B}$ near $T_{m}$ to approximately $2k_{\rm B}$ near the boiling temperature $T_{b}$ (Ref. Wallace (2002), p. 244). From the phonon theory, it can be interpreted that the degree of phonon freedom, $f$, per atom is effectively reduced from $f=6$ at $T_{m}$ to $4$ at $T_{b}$. Therefore, applying Eq. (5) to the liquid state leads to an overestimation of the phonon energy, resulting in the decrease in $E_{\rm st}$. The independent particle picture is not valid for the “phonons” of liquids. This invalidity of independent particle picture for liquids can be understood by considering the extreme case of an idea gas, where $C_{v}$ has the limiting value $f=3$, which is usually interpreted as the freedom of purely translational motions. When the frequency spectrum, $f_{v}(\omega)$, of atom velocity is calculated for an idea gas, we have a “phonon” spectrum, $g(\omega)$. Even though there is no restoring forces, atom collisions create the oscillatory behavior, yielding $f_{v}(\omega)$. Obviously, in this case, the phonon energy, Eq. (5) should not contribute to the $C_{v}$ of the gas. On this basis, it is apparent that applying Bose-Einstein statistics to “phonons” obtained from the velocity distribution of particles is inappropriate for fluids. The decomposition in Eq. (4) is not valid for liquids. Only the total energy $E_{\rm tot}$ has a physical reality.
III.3 Na

### v2

@@ 2.1.2 Construction of MD simulations
In this study, the internal energy $U$ must be evaluated over a wide range of temperature to cover from the solid to liquid states. To ensure that equilibrium was established at each $U$, an adiabatic MD run was performed for a fixed $E$ and repeated by changing the total energy. In adiabatic MD simulations, the control parameter is the initial temperature $T_{\rm in}$, at which the initial velocities of atoms are given through Eq. (7). In each single MD run, the simulation was continued until equilibrium was established. It is noted that there are two kinds of equilibrium conditions [69]. The one is the equilibrium condition with respect to temperature (thermal equilibrium). The equilibrium temperature is determined by
@@
Here, the brackets denote the particle average, whereas the bar indicates the time average. Thermal equilibrium is reached when the time average of $E_{\rm K}(t)$ converges. In the following, $T$ in all figures indicates this equilibrium temperature. The time required to reach this condition is referred to as the thermal relaxation time, $\tau_{T}$. For most of solids, $\tau_{T}$ is quite short, on the order of 0.1 ps. For liquids, $\tau_{T}$ becomes longer but is still less than 1 ps. Thus, in most cases the relaxation to equilibrium is controlled by the other condition, namely, the mechanical (structural) relaxation. The time required to this relaxation is referred to as the structural (mechanical) relaxation time, $\tau_{M}$. The time evolution of the particle-averaged displacements,
[…]
@@
is used to determine whether structural equilibrium is reached, where $t_{\rm sm}$ is the simulation time. For solids, it is considered an equilibrium state when $\langle\overline{\delta R_{j}(t)^{2}}\rangle_{j}$ is constant with respect to $t$. In contrast, for liquids, it is considered an equilibrium state when $\langle\overline{\delta R_{j}(t)^{2}}\rangle_{j}$ shows a linear dependence on $t$ over the entire simulation time $t_{\rm sm}$. For the latter case, the slope, $D=(1/6)\lim_{t\rightarrow\infty}\langle\overline{\delta R_{j}(t)^{2}}\rangle_{j}/t$, gives the diffusion coefficient $D$. For the liquids treated in this study, $\tau_{M}$ was on the order of a few ps. Examples of the time development of $E_{\rm K}$ and $\langle\overline{\delta R_{j}(t)^{2}}\rangle_{j}$ are provided in the Supplemental Material to Ref. [28].
2.2 Calculation conditions
@@
While several codes were used for the present FP-MD simulations, the majority of calculations were performed by Phase/0 [70]. Some were performed by a code written in-house (Osaka2k) or the VASP code [71]. All codes are pseudopotential methods using planewave expansion. The MD simulations employed time steps ranging from 0.72 to 1.2 fs and the total simulation time, $t_{\rm sm}$, varied from 2 to 10 ps, depending on the relaxation time. The volume of the crystal, $V$, was fixed at the experimental value, although normally the $V$ of the corresponding liquid phase is slightly larger than that of the crystal. The effect of volume change on $T_{m}$ is examined in subsections 3.3 and 3.4.
3 Results and discussion
3.1 Melting temperature
@@
Melting simulations for various materials were performed by FP-MD simulation by scanning a wide range temperature covering both the solid and liquid states. Table 1 summarizes the calculated $T_{m}$ together with experimental values for comparison. The potential type is indicated by NC (norm-conserving), US (ultrasoft), PAW (projector augmented waves) [72], along with the type of electron correlation energy functional (LDA or GGA). In MD simulations, melting occurs over a finite range of temperature, as shown later. Here this temperature range is referred to as the transition region. The calculated value of $T_{m}$ is then defined as the mid point of the transition region. The width, $W_{m}$, of the transition region is expected to approach zero as the size of supercell is increased. In the table, the deviation of the calculated value from the experimental value is indicated by the ratio, $r$, of the former to the latter. In all the cases studied, $r>1$ was obtained, meaning that the calculated values were always overestimated. It is our surprise to obtain large values of $r$ more than 2, despite the initial confidence in the DFT calculations. Large overestimations of $r>2$ are found when the material contains oxygen atoms. The involvement of oxygen for the large overestimation may be a reason why the coexistence method is widely used in the field of geophysics, because most materials important for that field are oxides [2, 3, 34, 73, 74, 36]. As2Se3 also has a large $r$. A recent study on c-BAs reports a large improvement in $T_{m}$ using the coexistence method [75]. The large overestimations of As2Se3 and of c-BAs could have the common origin. Only for Si, the calculated value is in a good agreement with experiment. Note that, in all the calculations in Table 1, the used volume of the unit cell is that of the crystal, which was obtained by structural optimization at $T=0$. Later, the influence of volume is investigated.
@@
In the table, the melting enthalpy $H_{m}$ is also listed in units of eV/atom. $H_{m}$ is obtained from the abrupt change in $E_{\rm st}$. In experiment, the change $H_{m}$ appears at a discrete temperature $T_{m}$, whereas melting occurred over a finite width $W_{m}$ in the present MD simulations. Hence, the accuracy of $H_{m}$ depends on the sharpness of the transition region. Small cell sizes give less reliable $H_{m}$ values. The accuracy of $H_{m}$ highly depends on the accuracy of $T_{m}$. These two quantities are related each other by
@@
where $\Delta S_{m}$ is the entropy of melting. This relationship says that an error of 0.1 eV in $H_{m}$ causes an error of 1000 K in $T_{m}$, provided $\Delta S_{m}/k_{\rm B}=1$. Because $\Delta S_{m}$ in many materials is of this order of magnitude, the accuracy of $H_{m}$ directly controls the accuracy of $T_{m}$. Again the calculation for Si shows a good agreement in $H_{m}$ with experiment. This confirms the internal consistency between calculated $T_{m}$ and $H_{m}$ for Si. The phonon energy, $E_{\rm ph}$, and thermal expansion energy, $E_{\rm te}$, can contribute $H_{m}$. However, the previous studies [27, 28] showed that these two contributions can generally be ignored, unless there is a large difference in thermal expansion between the solid and liquid phases. Because the melting temperature of most materials is high, the specific heat of the phonon contribution near $T_{m}$ becomes already the classical limit of the Dulong-Petit law, $C_{v}=3k_{\rm B}$, and thus there is no reason that an abrupt jump in $E_{\rm ph}$ occurs at $T_{m}$.
@@
$\alpha$-quartz 1700 4500 2.64 0.033 0.13 72 US+GGA 408 Phase
Glycerol 291 635 2.18 0.013 0.010 56 US+GGA 408 Phase
@@
Table 1: Comparison of calculated $T_{m}$ to the experimental value for various materials. $r$ indicates the ratio of the calculated $T_{m}$ to the experimental value. The melting enthalpy $H_{m}$ is also listed. Size indicates the size of the supercell used in MD simulations. The last block indicates the used potentials. $E_{\rm cut}$ for planewave expansion is given in eV units. $\Gamma$-point $k$ sampling is used in all the listed data. Experimental data are from [76] with the exception of the data for quartz [77].
@@
Several causes for the error in the calculated $T_{m}$ are investigated. One important cause is the use of periodic boundary conditions. An effect of the periodic boundary conditions is elimination of the surface effects. Because melting first occurs at the surface of a material, it can be argued that elimination of the surface makes melting less likely, thus increasing $T_{m}$. This effect could contribute to the overestimation of $T_{m}$. However, based on the good agreement between the calculated and experimental $T_{m}$ values for Si, this cannot be the primary reason for the large overestimation of $T_{m}$ unless a very large surface energy is involved. The introduction of defects could have a similar effect on melting. However, in our experience for quartz [28], the introduction of a defect in a supercell did not affect $T_{m}$ within the numerical accuracy.
@@
Another effect of the periodic boundary conditions is the creation of spurious reflections of atom movements, which yields an artificial barrier to such movements. For a given size of supercell, $L$, all long-wavelength phonons (those with wavelength $\lambda>L$) are eliminated. In a previous paper [28], by using a crude model, it is shown that this artificial barrier, $V_{0}$, is reduced only in a scale $1/L^{2}$, while the computational demands increase as $L^{3}$. The value of $V_{0}$ can be of the order of a few tens of eV with a cell size of $2\times 2\times 2$ for $\alpha$-quartz. However, to test this scaling, real calculations are needed by increasing the cell size. In practice, extending the size of supercells beyond those listed in Table 1 is very expensive. For a 512-atoms simulation, one month computation was required for a single temperature in our cluster machines (typically in an 8-core parallel configuration). Hence, the effect of supercell size can only be examined by reducing the supercell size. In the following, the effect of supercell size is examined by taking one example from each of the categories of semiconductors, metals, and oxide compounds. All the detailed data listed in Table 1 are seen in the Supplemental Materials.
3.2 Silicon
@@
The first example is Si. The cell sizes of $N=$8, 64, 216, and 512 atoms were examined. In Fig. 2, the structural energy, $E_{\rm st}$, and diffusion constant, $D$, are plotted as functions of $T$. Note that $E_{\rm st}$ is normalized by the number of atoms in the cell and is in units of K. This is convenient because the specific heat is immediately obtained in units of $k_{\rm B}$, making it easier to compare the obtained value to the classical limit $3k_{\rm B}$. This convention is used throughout this paper. In this figure, the data points are connected by lines to show the sequence of changing $T_{\rm in}$. As shown in Fig. 2, $E_{\rm st}$ is virtually constant against $T$ up to $T=1000$ K, indicating that the crystal potential is well described by the harmonic approximation. Finite values of $D$ appear at approximately $T=2000$ K, confirming the onset of melting. In accordance with the variation in $D$, $E_{\rm st}$ exhibits an abrupt increase there.
@@
Figure 2: The diffusion coefficient, $D$, and structural energy, $E_{\rm st}$, of Si calculated by NC+GGA. The cell sizes ($k$ meshes) are $N=8$ ($4^{3}$), 64 ($2^{3}$), 216 (1), and 512 (1). $E_{\rm cut}=204$ eV. The experimental value for $T_{m}$ is indicated by the dashed line. For each $N$, the energy origin is set to the ground state energy of that cell.
The figure shows that the $E_{\rm st}-T$ curve has the width, $W_{m}$, for the melting temperature. The $E_{\rm st}-T$ curve is expected to behave as a step function at $T_{m}$ and the magnitude of step must correspond to the latent heat. This width is the artifact due to use of finite-size cells, as clearly shown by its $N$ dependence. There is an unavoidable uncertainty in temperature, $\Delta T$, for finite-$N$ systems.
$\frac{\Delta T}{T}=\frac{\Delta E_{\rm K}}{\overline{E_{\rm K}}}=\frac{1}{\sqrt{N}}.$ (12)
@@
(Sometimes, it is referred to as the temperature fluctuation, but it may be better to term the kinetic-energy fluctuation [78, 79].) The heat exchange to the heat bath in MD simulations increases the kinetic-energy fluctuation, and this is another reason why the present method avoids the heat bath. This artificial width for melting is common in all materials, aside from the magnitude $W_{m}$. Furthermore, we notice in Fig. 2 that the transition curves, both $E_{\rm st}$ and $D$, exhibit sigmoidal behaviors, that is, negative slopes. The negative slope in $E_{\rm st}$ against $T$ means a negative specific heat, which is thermodynamically unstable [80, 81, 82]. The negative slope in $D$ against $T$ also means dynamical instability. This behavior was pointed out previously as the effect of finite-size cells [27, 28]. When the crystal melts, the energy distribution between the vibrational motions (in the solid portion) and translational motions (in the liquid portion) becomes unbalanced due to this energy fluctuation. At temperatures slightly below $T_{m}$, those atoms with sufficiently high kinetic energy (higher than $k_{\rm B}T_{m}$) begin to convert their vibrational motions to diffusing motions, decreasing the average kinetic energy $\langle E_{K}\rangle$ and increasing diffusion. In contrast, when this low-energy part of the distribution is overpopulated, the system temperature becomes lowered. This makes these low-energy diffusing atoms being trapped in the crystal potential, leading to a release of the energy in a form of emitting heat with a decrease in $D$. In this manner, an oscillatory behavior appears around $T_{m}$. In experiment, this sigmoidal behavior has been observed in clusters containing small numbers of atoms [83]. The transition region is narrowed as $N$ increases. On the thermodynamic limit, $N\rightarrow\infty$, $W_{m}\rightarrow 0$ is reasonably expected. However, it is noticeable that $W_{m}$ of 300 K remains even for the cell size of 512 atoms. The convergence of the energy fluctuation is slow.
The presence of this unstable region has some implications to the interpretation of MD simulations. When the temperature in a MD run is controlled by some means, such as thermostat, the system would happen to be trapped at either of the two extrema in the sigmoidal curve. If the system is relaxed to the maximum $T$ in the sigmoid, the obtained $T_{m}$ would become an overestimation. If the system is relaxed to the minimum $T$, then it would give an underestimation. These two extrema are spurious equilibria due to use of a finite $N$. The risk of this unstable region would be increased by using thermostats, because introduction of a thermostat usually increases the width of the energy fluctuation, even though the average temperature is well controlled. There is possibility that spurious equilibria happen in the Z method [37, 38]. Oscillatory behaviors between solid and liquid states are reported in [38].
@@
For Si, even though the convergence for the width $W_{m}$ is very slow, the calculated $T_{m}$, which is obtained by the mid point in the transition region, already agreed with the experimental value, when a cell size of $N=64$. Even a small cell of $N=8$ is not bad, aside from the $W_{m}$. From the viewpoint of construction of FP pseudopotential, Si is the easiest element [84, 85]. The potential is not deep and there are no semi-core levels. This may be the reason why the good agreement is obtained for $T_{m}$. In the literature, calculations of the melting temperature of silicon are reported by using thermodynamic integration [30, 86]. Although the agreement with the experiment is almost satisfactory, they gave an underestimation about 20 %. This underestimation could be accounted by the above-mentioned trapping at the spurious equilibrium with the lowest minimum. Because the thermodynamic integration method, many artificial steps of calculation, such as preparation of the reference state and adiabatic increase in the interaction parameter, are involved, it is difficult to analyze at which step the error occurs.
@@
For the liquid state, a decrease in $E_{\rm st}$ against $T$ can be seen in Fig. 2. This negative $T$ dependence is unphysical. In contrast, the total energy, $E_{\rm tot}$, is an increasing function of $T$, as should be. A full description of the $T$ dependence of $E_{\rm tot}$ will be given elsewhere [87]. Here, we only mention that this decreasing behavior of $E_{\rm st}$ is a consequence of applying Bose-Einstein statistics to the “phonons” of the liquid state. The decrease in $E_{\rm st}$ results in by subtracting overly large $E_{\rm ph}$ value from $\overline{E_{\rm gs}(t)}$ in Eq. (8). This temperature range ($T>T_{m}$) is already the classical regime, in which the Dulong-Petit law is expected to hold, if the phonon model is valid. All “phonon” modes are fully activated, which means that the thermal energy $(1/2)k_{\rm B}T$ is assigned to each mode $q$ of phonons according to Eq. (6), for which the independent particle description for phonons is assumed. In experiment, the $C_{v}$ of liquids decreases from approximately $3k_{\rm B}$ near $T_{m}$ to approximately $2k_{\rm B}$ near the boiling temperature $T_{b}$ (Ref. [53], p. 244). From the phonon theory, it can be interpreted that the degree of phonon freedom, $f$, per atom is effectively reduced from $f=6$ at $T_{m}$ to $4$ at $T_{b}$. Therefore, applying Eq. (6) to the liquid state leads to an overestimation of the phonon energy, resulting in the decrease in $E_{\rm st}$. The independent particle description is not valid for the “phonons” of liquids. This invalidity of the independent particle description for liquids can be understood by considering the extreme case of an ideal gas, where $C_{v}$ has the limiting value $f=3$, which is usually interpreted as the freedom of purely translational motions. When the frequency spectrum, $f_{v}(\omega)$, of atom velocity is calculated for an ideal gas, we have a “phonon” spectrum, $g(\omega)$. Even though there is no restoring forces, atom collisions create the oscillatory behavior, yielding $f_{v}(\omega)$. Obviously, in this case, the phonon energy, Eq. (6) should not contribute to the $C_{v}$ of the gas. On this basis, it is apparent that applying Bose-Einstein statistics to “phonons” obtained from the velocity distribution of particles is inappropriate for fluids. The decomposition in Eq. (5) is not valid for liquids. Only the total energy $E_{\rm tot}$ has a physical reality.
3.3 Sodium